OpenStudy (yueyue):

Let f(x)=(sin(4x))/x and g(x)=3x^2+2. Find the limit of f(x) + g(x) as x approaches 0.

OpenStudy (yueyue):

The answer cannot be undefined.

OpenStudy (anonymous):

To begin with, the limit of f(x) + g(x) as x approaches 0 equals the limit of f(x) as x approaches 0 plus the limit of g(x) as x approaches 0. The limit for g(x) is obviously 2, so we can set that aside (but remember to add it at the end).

OpenStudy (anonymous):

multiply f(x) by 4/4 so it becomes \[4*\frac{ \sin4x }{ 4x } and we know \frac{ sinx }{ x} =1 at x \rightarrow0\]

OpenStudy (anonymous):

ang g(x) will be 2 so ans is 4+2 = 6

OpenStudy (yueyue):

Alright, got it. Thank you!

OpenStudy (anonymous):

no prb..:D